OpenStudy (anonymous):

EASY QUESTION!!!!! A test consists of multiple choice questions worth 3 points and essay questions worth 5 points. The test has 17 questions that are worth a total of 65 points. Use a model to determine how many multiple choice questions are on the test. A. 10 B. 11 C. 12 D. 13

OpenStudy (bahrom7893):

3x + 5y = 65 x+y = 17 solve for x

OpenStudy (anonymous):

Is it 10?

OpenStudy (bahrom7893):

x = 17-y 3(17-y) +5y = 65 51-3y +5y = 65 51 + 2y = 65 2y = 65-51 = 14 y = 7 x = 17-7 = 10, YES

OpenStudy (anonymous):

He he I knew it. Thanks :-)